300g of an unknown metal at 200°C was placed in a 50g Aluminium pot (C =4.184J/gC ) at 21°C.The final temperature of the mixture was 29°C. What is the specific heat capacity of the metal?

1
Expert's answer
2022-01-10T13:57:52-0500

Heat transfer = m c ΔT


Heat released by the metal = (300 g) × (c J/g°C) × [(29 - 21)°C] = 2400c J

Heat gained by the water = (100 g) × (4.184 J/g°C) × [(29 - 21)°C] = 3347.2 J

Heat gained by the pot = (50 g) × (0.90 J/g°C) × [(29 - 21)°C] = 360 J


Heat released by the metal = (Heat gained by the water) + (Heat gained by the pot)

2400c = 3347.2 + 360

Specific heat capacity of the metal, c = 1.54 J/g°C